Could you please help me with this problem, I need to add 1/3 + (-5/6) + (-1/2). Please help me with this. Thank you.
Answer by rfer(16322)

1/3+(-5/6)+(-1/2)
1/3-5/6-1/2
2/6-5/6-3/6=-6/6=-1